Rosita's has a cost of equity of 13.76 percent and a pretax cost of debt of...

  1. Rosita's has a cost of equity of 13.76 percent and a pretax cost of debt of 8.5 percent. The debt-equity ratio is .60 and the tax rate is 34 percent.
    1. What is Rosita’s WACC if his debt to equity ratio is 0.3?
Answer #1
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
13.76 = Unlevered cost of equity+0.6*(Unlevered cost of equity-8.5)*(1-0.34)
Unlevered cost of equity = 12.27
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 12.27+0.3*(12.27-8.5)*(1-0.34)
Levered cost of equity = 13.02
D/A = D/(E+D)
D/A = 0.3/(1+0.3)
=0.2308
Weight of equity = 1-D/A
Weight of equity = 1-0.2308
W(E)=0.7692
Weight of debt = D/A
Weight of debt = 0.2308
W(D)=0.2308
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 8.5*(1-0.34)
= 5.61
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.61*0.2308+13.02*0.7692
WACC =11.31%